OpenStudy (anonymous):

4x^4-12x^2+5=0 solve

OpenStudy (dumbcow):

factor (2x^2-5)(2x^2-1) = 0 2x^2 -5 =0 x^2 = 5/2 x = +-sqrt(5/2) 2x^2 -1 = 0 x^2 = 1/2 x = +-sqrt(1/2)